OpenStudy (anonymous):

can someone check these for me? This is based out of a standard deck of 52 cards. · a face card given the card is a not a club. · a heart given the card is black. · a Jack, given that the card is red. my answers: 1. 9 out of 12 face cards not being a club, 9/12 = 3/4 2. cutting the deck in half, 12/26 reduces to 6/13 3. 1/26 chance of getting a red Jack

OpenStudy (anonymous):

I forgot to say, this is out of probability

Directrix (directrix):

1)Question: (a face card given the card is a not a club) --> does this mean that the card drawn cannot be any type club or just not a "face" club? 2) P (a heart given the card is black) --> I don't think there are any black hearts in the deck. My working answer there is P(Black Heart) = 0/52 = 0. 3) P( red Jack) = 2/52 = 1/26. [There are 2 red Jacks in the deck, Jack of Diamonds and Jack of Hearts.]

OpenStudy (anonymous):

ooops, i think i messed the whole question up, give me a sec

Directrix (directrix):

There are 13 clubs in the deck. I am asking if " (a face card given the card is a not a club) should say ( a face card given the card is not a face club card) or if the question is asking (a face card given the card is a not a club of any type.)

OpenStudy (anonymous):

Any kind of face card given it isn't a club, my mistake.

OpenStudy (anonymous):

"(a face card given the card is a not a club of any type.)" thats correct

Directrix (directrix):

P(Any kind of face card given it isn't a club) = 9/52. What do you think?

OpenStudy (anonymous):

That seems correct to me. Non reducible but it can't apply if that is the right answer. I appreciate your assistance!